Find the first partial derivative of



s(y, z) = z ^ 2 * tan yz

1
Expert's answer
2021-12-07T09:47:48-0500
"\dfrac{\partial s}{\partial y}=\dfrac{z^3}{\cos^2(yz)}"

"\dfrac{\partial s}{\partial z}=2z\tan(yz)+\dfrac{yz^2}{\cos^2(yz)}"